OpenStudy (anonymous):

Solve V = π r2h for h.

OpenStudy (whpalmer4):

\[V = \pi r^2 h\]You want \(h\) all by itself, so divide both sides of the equation by everything that is multiplied together with \(h\) in the current equation.

OpenStudy (texaschic101):

V = (pi)r2h for h V/(pi)r2 = h

OpenStudy (whpalmer4):

That's a somewhat sloppy way of writing it: how do we know for sure that you don't mean h = (V/pi)r^2 ? h = V/(pi r^2) eliminates any possible ambiguity.

OpenStudy (texaschic101):

your right......that could be confusing....good catch

OpenStudy (whpalmer4):

(or typesetting the equation as \[h = \frac{V}{\pi r^2}\] eliminates the ambiguity as well)

OpenStudy (whpalmer4):

no point in getting the right answer if the person on the other end can still interpret it as a wrong answer :-)

OpenStudy (texaschic101):

correct again...you should be a teacher :)

OpenStudy (whpalmer4):

this discussion comes up frequently when looking at my kid's homework — many correct answers which simply can't be deciphered by anyone but him :-)